Find two integers whose sum is -9 and product is 20
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Find two integers whose sum is -9 and product is 20
===========
Try pairs of factors of 20.

Answer by stanbon(75887)

Find two integers whose sum is -9 and product is 20
----
x + y = -9
x*y = 20
-----
Substitue for "x" and solve for "y"::
(-9-y)y = 20
-------
-y^2 - 9y - 20 = 0
----
y^2 + 9y + 20 = 0
(y+4)(y+5) = 0
y = -4
x = -5
-------------
Cheers,
Stan H.